Ticket Number: HEA-NOC/20060412-1 Ticket Status: UPDATE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20060412 17:14 UTC+1 Problem Start: 20060412 16:46 UTC+1
Ticket Update: 20060705 17:40 UTC+1 Problem End:
Co-ordinate the relocation of services and servers to servecentric data
20060412 16:49 SY Created a new list with co-location client contacts to
inform and co-ordinate the best times for their services to be relocated.
([log in to unmask]). Notified clients of the requirement
to relocate and to arrange meeting in advance to co-ordinate and arrange
suitable dates for the move for minimium disruption to their services &
20060426 16:03 SY Meeting Frank O'Brian in BH on Wed 3rd May to discuss
his equipment move and a possible visit for the EPA to servecentric.
20060524 12:04 SY IRIS move scheduled for Friday Morning 26 May,
notification added to IRIS web home page.
HEA will relocate their webserver HEAweb and Oracle backend on Saturday
27th, they will notify there users.
Qualifax and ILUG servers also moving on Sat 27th in Morning 9:30am to
20060526 14:33 SY IRIS (koranov) service relocated physically to new data
centre, services restored successfully, ticket open for monitoring.
20060529 11:05 SY CosmoGrid breezlit windows server to be relocated Monday
11:30am to 3:00pm. EPA equipment and service move 11:00 to 3:00pm, TSSG
servers move Monday 11:00am to 3:00pm. IRCHSET and UCD-RIS server move
scheduled for Thursday 1st June.
20060529 11:27 SY John Allman from DIAS has arranged to come in on
Thursday 1st June to oversee to relocation of the DIAS cosmogrid server,
so the service will not be offline today Monday 29th May.
20060529 11:46 SY ESAT co-located morbo server scheduled to be moved on
Thursday morning 1st June, copying ticket to ESAT BT noc.
20060529 12:36 SY TSSG servers relocated today Mon 29th May, 12:30 to
20060531 14:19 SY IRCHSET E-GRANTS scheduling, June 1st comprehensive
labelling by kerna consultants. June 7th kerna dissassemble and pack up
servers and services, June 8th kerna reassemble servers and services in
servecentric and test fully.
20060607 10:47 SY TSSG final two servers scheduled to move on Sat 17th
20060607 16:20 SY BT Ireland/ESAT co-located server morbo.esat.net to be
relocated to HEAnet new data centre on Thursday morning, 8th June from
10:00am to 12:30am. ESAT to be notified when the server is relocated and
back on-line for testing.
20060616 12:39 SY TSSG WIT server relocated successfully, IRCSET office
servers moved from co-location room to HEA area. ILUG to delete and
removal arranged for their old HP server, AC units in co-lo room switched
off including condensor in car park. Co-location services and servers move
completed. Only remains to decommission racks in BH room.
20060705 17:39 SY Moving backup server(orion), it's disk array and L20
tape library tommorrow 6th July to servecentric data centre.
This ticket can be monitored at http://www.hea.net/tickets/20060412-1